OpenStudy (anonymous):

Solve, using the quadratic formula: 6x^2 = -5x + 25 (Points : 1) {-5/3, 5/2} {5/3, -5/2} No real solutions

OpenStudy (nathan917):

@miamia6547 Simplifying 6x2 = -5x + 25 Reorder the terms: 6x2 = 25 + -5x Solving 6x2 = 25 + -5x Solving for variable 'x'. Reorder the terms: -25 + 5x + 6x2 = 25 + -5x + -25 + 5x Reorder the terms: -25 + 5x + 6x2 = 25 + -25 + -5x + 5x Combine like terms: 25 + -25 = 0 -25 + 5x + 6x2 = 0 + -5x + 5x -25 + 5x + 6x2 = -5x + 5x Combine like terms: -5x + 5x = 0 -25 + 5x + 6x2 = 0 Factor a trinomial. (-5 + -2x)(5 + -3x) = 0 Subproblem 1 Set the factor '(-5 + -2x)' equal to zero and attempt to solve: Simplifying -5 + -2x = 0 Solving -5 + -2x = 0 Move all terms containing x to the left, all other terms to the right. Add '5' to each side of the equation. -5 + 5 + -2x = 0 + 5 Combine like terms: -5 + 5 = 0 0 + -2x = 0 + 5 -2x = 0 + 5 Combine like terms: 0 + 5 = 5 -2x = 5 Divide each side by '-2'. x = -2.5 Simplifying x = -2.5 Subproblem 2 Set the factor '(5 + -3x)' equal to zero and attempt to solve: Simplifying 5 + -3x = 0 Solving 5 + -3x = 0 Move all terms containing x to the left, all other terms to the right. Add '-5' to each side of the equation. 5 + -5 + -3x = 0 + -5 Combine like terms: 5 + -5 = 0 0 + -3x = 0 + -5 -3x = 0 + -5 Combine like terms: 0 + -5 = -5 -3x = -5 Divide each side by '-3'. x = 1.666666667 Simplifying x = 1.666666667 Solution x = {-2.5, 1.666666667}